(a) A bowl of water has a temperature of 50◦C. It is put into a refrigerator where the temper-
ature is 5◦C. After 0.5 hours, the water is stirred and its temperature is measured to be
20◦C. It is then left to cool further, take t in Newton’s law of cooling to be measured in
minutes. Use Newton’s law of cooling to predict when the temperature will be 10◦C.

(b) Verify that the given family of functions solves the differential equation.
(i) dy
dt = (1 − 2t)y
2
, y =
1
C − t + t
2
.
(ii) dy
dt = y
2
sin t, y =
1
C + cost
.
(c) Evaluate the integral Z 1
0
x
2
(

4 − x
2)
3
dx
1
Expert's answer
2020-07-02T19:27:28-0400

(a) Equation of newton's law of cooling,

"\int \frac{dT}{(T-T_0)} = - \int k dt"


Where T is temperature at any time t and T0 is the temperature of surrounding.

Given, (i) at t = 0, T = 500C = 323K

(ii) at t = 0.5 hrs, T = 20 C = 293 K

Temperature of the surrounding is T0 = 5 C = 293 K


Solving equation we obtain,

"ln(T-T_0) = -kt + lnC"

applying condition (i)

"lnC = ln (45)"

applying (ii) condition,

"k = 2ln(3)"

Hence final equation will be,

"ln(T-293) = -2ln(3) t + ln(45)"

If T = 10 C = 283 K, then for time calculation,

"ln(283-278) = -2ln(3)t + ln(45) \implies t = 1 hour"


b (i) Given "\frac{dy}{dt} = (1-2t)y^2" and "y = \frac{1}{c-t+t^2}"

Differentiating y with respect to x, 

"\frac{dy}{dx} = -(\frac{1}{c-t+t^2})^2(-1+2t) = (\frac{1}{c-t+t^2})^2(1-2t)"

replacing value of y 

"\frac{dy}{dt} = (1-2t)y^2"


(ii) "\frac{dy}{dt} = y^2sin(t)" and "y = \frac{1}{c+cost}"

differentiating both sides with respect to x,

"\frac{dy}{dt} = -(\frac{1}{c+cost})^2(-sint) = (\frac{1}{c+cost})^2(sint) = y^2sin(t)"


(iii) "\int_0^1 x^2 (\sqrt{4-x^2})^3dx"

let "x = 2sin(u)" then "dx = 2cos(u) du"

Let us solve integration first without limit

"I=\int 8cos(u)sin^2(u)(4-4sin^2(u))^{3/2}du
=\int 8cos(u)sin^2(u)(4cos^2(u))^{3/2}du =\int 64cos^4(u)sin^2(u)du = \int 64cos^6(u)(1-cos^2(u))du"


Integrating it,

"I = -\frac{x(4-x^2)^{5/2}}{6} + \frac{x(4-x^2)^{3/2}}{6}+x(4-x^2)^{1/2} + 4asin(\frac{x}{2})"

putting limits,

"\int_0^1 x^2 (\sqrt{4-x^2})^3dx = \frac{2\pi}{3}"